OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 4/5, what is the equation of the line? A. 4y - 5x = 5 B. Y - 4x = 5 C. 5y - 4x = 5

OpenStudy (anonymous):

y = mx + b m = 4/5, the point (0, 1) is the y-intercept, so b = 1 so, y = (4/5)x + 1

OpenStudy (mokeira):

\[\frac{ 4 }{ 5 }=\frac{ y-1 }{ x-0 }\] cross multiply \[4(x-0)=5(y-1)\] \[4x=5y-5\] \[5y=4x+5\] or \[4x-5y=5\]

OpenStudy (mokeira):

sorry it is suposed to be 5y-4x=5